OpenStudy (anonymous):

Solve the equation and write the values of the variable that make the denominator zero. (3/x)=((15/8x)+3)

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

well, what do we have in the denominators?

OpenStudy (anonymous):

x and 8x and 1. so im thinking multiply 15/8x by just 1, and then do i multiply 3 or 3/x by 8?

OpenStudy (amistre64):

just multiply thru the whole thing by 8x to solve for x and the zeros for the denominators are when x or 8x equal zero

OpenStudy (amistre64):

8x [(3/x)=((15/8x)+3)] 3(8)=15+3(8x)

OpenStudy (anonymous):

im confused okay now i have ((24x/8x) = (144x/72x)) should i divide it all by 8 or 8x now???

OpenStudy (anonymous):

Whats the problem ? Start again ?? May I answer you ???

OpenStudy (anonymous):

Yeah sure. The problem is (3/x)=((15/8x)+3) and i've multiplied it all by 8x so now I have ((24x/8x)=(120x/64x) + (24x/8x))

OpenStudy (anonymous):

I want to know what I should do from this point

OpenStudy (anonymous):

Look : If we have : 3/x=15/8x+3 *8x we have : 24x/x=15+24x and have : 23x=15+24x I think it s right :)

OpenStudy (anonymous):

Could you explain why you did that? I know my end product is going to "x=(3/8)". Will what you gave me produce that?

OpenStudy (anonymous):

OK ! Look : 8x * 3/x = ?

OpenStudy (anonymous):

24x/8x

OpenStudy (anonymous):

No ! Your fall was in there :) we multiplied 8x just in 3 not in x :) x is Denominator there :) Got it ? Look : 2 * 1/2 = 2*1/2 not 2*1/2*2 OK ?:)

OpenStudy (anonymous):

@rrey :Got it ?:)

OpenStudy (anonymous):

Why wouldn't I multiply the denominator? If I do something to the top I have to do it to the bottom too, don't I?

OpenStudy (anonymous):

I said for u look if we do that think u said have : 2 * 1/2 = 2*1/2 not 2*1/2*2 Got ?:)

OpenStudy (anonymous):

@rrey : ????

OpenStudy (anonymous):

Oh right.

OpenStudy (anonymous):

Got it ?:)

OpenStudy (anonymous):

Yeah. Thank you.

OpenStudy (anonymous):

And dont have any problem ?:)